A fly wheel has a moment of inertia of 3.8kgm^2.what constant is required to increase its frequency from 2rev/sec to 5rev/sec in 6 revolutions?
1
Expert's answer
2020-05-12T09:32:46-0400

6 revolutions correspond to angular displacement of "\phi=6\cdot2\pi=12\pi".

To increase the angular velocity, there must be angular acceleration:


"\alpha=\frac{\omega_2^2-\omega_1^2}{2\phi}=2\pi^2\cdot\frac{n_2^2-n_1^2}{\phi}=11\text{ rad/s}^2."

The torque required to provide such acceleration is


"\tau=I\alpha=41.8\text{ N}\cdot\text{m}."